### v1

# Ludax: A GPU-Accelerated Domain Specific Language for Board Games

Abstract
Games have long been used as benchmarks and testing environments for research in artificial intelligence. A key step in supporting this research was the development of game description languages: frameworks that compile domain-specific code into playable and simulatable game environments, allowing researchers to generalize their algorithms and approaches across multiple games without having to manually implement each one. More recently, progress in reinforcement learning (RL) has been largely driven by advances in hardware acceleration. Libraries like JAX allow practitioners to take full advantage of cutting-edge computing hardware, often speeding up training and testing by orders of magnitude. Here, we present a synthesis of these strands of research: a domain-specific language for board games which automatically compiles into hardware-accelerated code. Our framework, Ludax, combines the generality of game description languages with the speed of modern parallel processing hardware and is designed to fit neatly into existing deep learning pipelines. We envision Ludax as a tool to help accelerate games research generally, from RL to cognitive science, by enabling rapid simulation and providing a flexible representation scheme. We present a detailed breakdown of Ludax’s description language and technical notes on the compilation process, along with speed benchmarking and a demonstration of training RL agents. The Ludax framework, along with implementations of existing board games, is open-source and freely available.
1 Introduction
For the past 75 years, games have served as vital tests and benchmarks for artificial intelligence research. While many specific games have been completely solved schaeffer2007checkers or optimized beyond the abilities of the strongest human players campbell2002deep ; silver2017mastering , the general space of games remains a fertile ground for measuring improvements in reasoning, planning, and strategic thinking. A critical part of this progress, however, is the ability to test approaches and algorithms on a set of environments that are both diverse and computationally efficient.
To help drive further games and learning research, we introduce Ludax: a domain-specific language for board games that compiles into GPU-accelerated code written in the JAX library jax2018github . Ludax draws on two main inspirations: (1) Ludii Piette_2020_Ludii , a general purpose description language for board games capable of representing more than 1400 games from throughout history and around the world, and (2) PGX koyamada2023pgx , a collection of optimized JAX-native implementations of classic board games and video games designed to facilitate rapid training and evaluation of modern reinforcement learning (RL) agents. Ludax presents a flexible and general-purpose game representation format that can be leveraged for efficient simulation and learning on modern computing hardware.
Ludax currently supports two-player, perfect-information, turn-based board games played by placing pieces onto empty board cells. While this is a narrow class of games when compared to the full breadth of human designs and the set of mechanics implemented by Ludii, it is still broad enough to capture a wide range of existing games (e.g. Connect Four, Pente, Hex, …) as well as many unexplored novel games and variants that fall within that class. Further, Ludax is designed to be easily expandable – like with Ludii, implementing new game mechanics in Ludax only requires implementing new atomic components in the underlying description language. These components can then be combined compositionally with existing elements of the language to produce an entirely new range of possible games, instead of each game needing to be implemented separately.
Another design goal for Ludax is ease of use, both in terms of game design and experimentation. The syntax of the description language is “ludemic” Piette_2020_Ludii – splitting game rules into clear sections governing the game’s setup, play mechanics, and end conditions. Like with Ludii, game programs in Ludax resemble English descriptions of rules (see Figure 1). Further, by leveraging the structure of the existing PGX library, environments instantiated in Ludax can be easily combined with existing frameworks for GPU-accelerated search, reinforcement learning, or evolution deepmind2020jax ; tang2022evojax . Ludax also supports a basic web interface for interactive debugging, with the aim to provide support for online interactive experiments in the future.
To our knowledge, Ludax is the first game description language which compiles into GPU-accelerated code. In the following sections, we provide a detailed description of the language syntax, compilation process, and Ludax’s expressive range. We also provide speed benchmarking compared to both Ludii and PGX, as well as an initial demonstration of training learned agents. Finally, we conclude with a discussion of potential use cases and future directions. Ludax is open-source under the Apache 2.0 license, and the code is available at: https://github.com/gdrtodd/ludax.
Figure 1: Example game descriptions for Reversi, Pente, and Yavalax in the Ludax game description language, along with literal translations into natural language.
2 Related Work
While our core contribution of combining game description languages (GDLs) with hardware accelerated learning environments is novel, there is extensive research in each of these domains separately.
Game Description Languages: Game description languages have been used for many years and in a variety of domains. The Stanford GDL Love_2008_GDL ; Genesereth_2014_GGP ; Schiffel_2014_Representing ; Thielscher_2017_GDLIII is among the most influential, helping to popularize research in general game playing Pitrat_1968_GGP through its use in the International General Game Playing Competition Genesereth_2005_GGP ; Genesereth_2013_GGP . Other notable examples include VGDL ebner2013towards ; Schaul_2013_VGDL ; Schaul_2014_Extensible (primarily known from its use in the General Video Game AI framework Perez_2019_GVGAI ), RBG Kowalski_2019_Regular , Ludi Browne_2009_PhD , and its successor Ludii Piette_2020_Ludii . GDLs have also been used to describe the rules of card games font2013card as well as to represent human goals in naturalistic simulated environments davidson2022creativity ; davidson2025goals . Modern game description languages have tended to move away from a basis in formal logic in favor of greater human usability, though there are benefits in efficiency gained by the use of regular languages Kowalski_2020_Efficient .
GPU-Accelerated Environments: Recent years have seen a proliferation of learning environments implemented in the JAX library or other frameworks that enable hardware (typically GPU) acceleration. Examples include single-agent and multi-agent physics simulators freeman1brax ; Makoviychuk_2021_Isaac ; Bettini_2022_VMAS , ports of both classic and recent reinforcement learning tasks Dalton_2020_Accelerating ; gymnax2022github ; koyamada2023pgx ; matthews2024craftax , combinatorial optimization problems bonnet2023jumanji , multi-agent coordination problems rutherford2023jaxmarl , and driving simulators Gulino_2023_Waymax ; kazemkhani2025gpudrive . While these efforts have spurred significant progress and span a wide range of domains and task formulations, each of them implement a fixed environment or set of environments. As such, they cannot easily be extended to novel environments without first writing new hardware-accelerated code.
3 Description Language Details
Ludax’s game description language draws heavily on the Ludii description language, particularly in its use of “ludemic” syntax that represents game rules in terms of high-level and easily-understandable components Piette_2020_Ludii . The complete grammar file and syntax details are available in the Supplemental Material.
3.1 Equipment
The equipment section contains information about the physical components used by the game. Currently, this only specifies the size and shape of the board (i.e. whether it is square, rectangular, hexagonal, or hexagonal-rectangular). The dimensions and shape of the board are used during compilation to help pre-compute certain game-relevant properties, such as the board indices corresponding to lines of specific lengths. In future versions of Ludax, the equipment section will also detail the different pieces used by each player if the game specifies more than one.
3.2 Start Rules
The start section is an optional section that contains the rules for the game’s setup. For most games, play begins on an empty board and the start section is omitted. In some games, such as Reversi (see Figure 1, left panel), pieces are placed in a particular arrangement at the start of play.
3.3 Play Rules
Typically, the play rules of each game are the most involved, as they detail the core mechanics and dynamics of the game. The play section is itself broken into one or more subsections called “play phases.” Each phase has its own rules for player actions and turn-taking, as well as specific conditions for when to transition to another phase. Most games have only a single phase in which players alternate turns until the game is over, specified with the repeat keyword. Some games include a once-through phase that progresses through the turn order a single time before advancing to the next phase. The sequence of player turns is specified independently for each phase. For instance, Yavalax (Figure 1, right panel) begins with the first player making a single move (i.e. (once-through (P1) ...) ) before both players alternate taking two turns for the rest of the game (i.e. (repeat (P2 P2 P1 P1) ...) ).
The core of each phase is a “play mechanic” that encodes the ways that players take their turns. In the context of reinforcement learning, a play mechanic specifies both the action space ($\mathcal{A}$) and the transition function ($\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S})$. At a lower level, each play mechanic also defines a “legal action mask function” that returns whether each action is valid from the current game state. Currently, Ludax supports only one kind of play mechanic: place. A place mechanic’s primary argument is a destination constraint which specifies where a piece may be placed on a given player’s turn. For many games, such as Tic-Tac-Toe, this is simply the set of empty board positions. For some games, however, the destination constraint is more involved: in Connect Four, legal actions are empty spaces that are on the bottom edge of the board or immediately above an occupied position (see subsection 3.5 for a discussion of how actions are represented more generally in Ludax). Even further, some games have what we call result constraints which require that a legal action results or doesn’t result in a board with specific properties. Yavalax and Reversi both use result constraints: the former forbids players from placing a piece that forms a line of five or that forms only a single line of four, whereas the latter requires players to place a piece in a way that “sandwiches” one or more of their opponent’s pieces in a line. Finally, a play mechanic may optionally specify one or more effects that modify the game state after the action is performed. Effects are used to handle mechanics like capturing or flipping pieces, as well as updating each player’s score (if the game uses score). Both Reversi and Pente use play effects to handle flipping and capturing pieces, respectively, with Pente also using the score as an alternate winning condition.
Throughout this section, we have been referring to various properties of a game state and relationships between pieces / positions (e.g. whether pieces are “sandwiched,” whether a line is formed, whether a piece is adjacent to another, …). These are the lowest-level component’s of Ludax’s description language and are referred to collectively as masks, functions, and predicates. A mask takes in the current game state and returns a boolean array over each position on the board. Some masks, like occupied or edge, take additional grammatical arguments which might specify a particular player or region of the board. [^1] The adjacent mask is a special case – it takes another mask as an additional argument and returns the board positions adjacent to any of the active positions in the original mask. A function similarly takes in the current game state and returns a single non-negative integer. In Ludax’s current form, line is probably the most commonly-used function – it returns the number of contiguous lines of a given player’s pieces on the board, with a specified length and orientation. Lastly, a predicate maps from a game state to a single boolean truth value. Many predicates operate over the outputs of masks and functions, such as exists or equals, though some like mover_is are computed directly from game states. Crucially, the outputs of masks, functions, and predicates can be combined compositionally using first-order logic (excluding quantification) to form more complicated expressions. So, the condition “if Player 2 makes a line of 4 in a row or a diagonal line of 3…” would be rendered as follows:
(and (mover_is P2) (or (line 4) (line 3 orientation:diagonal)))
Note that, for ease of use, Ludax automatically interprets the presence of a bare function inside a boolean operator as indicating a non-zero value. So, (line 4) is equivalent to (>= (line 4) 1).
3.4 End Rules
The last section of a game description in Ludax details the criteria that terminate a game. The end section contains one or more “end conditions” – these are applied in order, with the first condition to activate determining the ending behavior (i.e. which player wins or if the game ends in a draw). If none of the conditions activate, then the game continues. For instance, Tic-Tac-Toe includes both the end conditions (if (line 3) (mover win)) and (if (full_board) (draw)), with the draw condition only triggering if the “three in a row condition” is not met. End conditions also frequently refer to a player’s score, which is updated or set as a result of an action’s effects (see above).
3.5 Design Considerations
While Ludax draws heavily from the Ludii description language, there are some important differences which go beyond just changes in syntax. The first of these relates to how both systems represent a game’s action space. One of the design goals of Ludii is that game descriptions should resemble as much as possible the rules in natural language. In Connect Four, for instance, players take a move by dropping a piece into one of seven columns of the board, at which point the piece falls until it reaches the bottom or rests on another piece. Accordingly, the canonical representation of Connect Four in Ludii features pieces that “Drop” into the “LastColumn” chosen by the player (PGX implicitly represents the game in a similar way). As mentioned above, however, Ludax represents the action space differently: players simply place a piece onto an empty board cell, with actions that are not directly above an existing piece or the bottom of the board marked as illegal. Mechanically, the two implementations of Connect Four are identical – the difference lies in how they are encoded (especially to simulated players or reinforcement learning agents). The “column-based” representation has many advantages (it matches the physical properties of the game in real life and lowers the branching factor), but it is also game-specific. While Ludax also strives to represent game descriptions intuitively, we primarily aim to provide a unified representation format across games, such that general game-playing agents can more easily transfer knowledge and expertise from one game to another. As such, the size and form of the action space for any place-based game is determined only by the size and shape of the board. This choice is also partially motivated by the specifics of working with the JAX library (see Section 4) and has implications for benchmarking and downstream use-cases (see Section 6).
4 Compiling Game Descriptions into Game Environments
In this section, we describe the high-level approach used to map from programs in the Ludax game description language to hardware-accelerated simulation environments. While Ludax specifically instantiates board game environments using the Lark Python library, the general approach is flexible enough to be used with different domains and parsing toolkits. Broadly speaking, Ludax operates by defining the leaves of the grammatical parse tree (i.e. individual masks, functions, and predicates) as atomic functions written in JAX, which are then dynamically composed from the bottom-up to form higher-level operators used by the environment class. Consider again the following game expression:
(and (mover_is P2) (or (line 4) (line 3 orientation:diagonal)))
During compilation, the leaf-level nodes (i.e. (mover_is P2) and (line 4)) are converted into JAX functions which map from the current game state to (in this case) a boolean truth value, and those functions are then passed up the parse tree. Higher-level nodes, such as (and ...), receive the JAX functions corresponding to each of their children and return a new JAX function that also takes the game state as input and implements the appropriate operation (in this case, boolean conjunction). In pseudocode, using the Lark library’s Transformer paradigm, this looks like the following:
def predicate_and(self, children):
def predicate_fn(state):
children_values = [child_fn(state) for child_fn in children]
return all(children_values)
return predicate_fn
In actuality, both the “children functions” and the combined “predicate function” must be written to be compatible with JAX’s vectorization scheme and just-in-time (JIT) compilation. This imposes a number of implementational constraints, most notably that the size and shape of all arrays must be fixed at compile time. This means, for instance, that the dimensions of the “legal action mask” (and, hence, the size of the action space in general) cannot change as the game progresses. In addition, values like the number of iterations in a loop or the positions of a lookup mask must essentially be “pre-specified.” Crucially, however, values that are determined during parsing (such as the number of children for a given node, or the value of any arguments) can be safely passed into compiled JAX functions as static constants. This fact is what allows Ludax to create JAX functions dynamically that nonetheless obey the constraints of vectorization and JIT compilation. At the top of the parse tree, these composed JAX functions are ultimately used to define the behaviors that appear in the environment’s step function, such as applying the player’s action to the board and handling move effects.
We next discuss some of the specific optimizations used by Ludax. In general, these are not global optimizations: they apply only to certain compositions of game rules and mechanics. Our approach is to deploy these optimizations when they are available and to “fall back” on slower but more general solutions when they are not.
Precomputation: An important optimization used by the PGX library (and JAX environments more generally) is to express functions as batched matrix operations rather than iterative procedures. For instance, rather than checking for a line of pieces in Tic-Tac-Toe by starting at the position of the last move and scanning out in each direction (as Ludii’s implementation does), PGX hard-codes the set of board indices that correspond to each possible line of three in the game (i.e. [[0, 1, 2], [0, 3, 6], [0, 4, 7], ...]) and performs a single multi-dimensional index into the board array – if any of the of the board index triples all correspond to positions occupied by a single player, then the game is over. Ludax adopts and generalizes this approach: during parsing of line, for example, the line indices are computed with respect to the size and shape of the game board (i.e. rectangular, hexagonal, …) as well as the length and orientation of the desired line (i.e. diagonal, vertical, …). Again, because these values depend only on attributes that are determined during parsing, they can be passed into JAX functions as constants. Precomputation naturally causes a trade-off between compile-time and run-time efficiency. In our case, we opt to use precomputation whenever possible, though some masks and functions cannot be expressed this way.
Dynamic State Attributes: Different games require tracking different kinds of information about the current game state. Most obviously, some games track a score for each player while others do not. When Ludax compiles a game, it automatically extracts the attributes required to instantiate a game state and omits the others, thereby reducing the memory footprint of the entire state object. More importantly, Ludax also automatically adds intermediary computations to each call of the environment’s step function that help speed up later mask, function, or predicate evaluations. For example, in Hex, the game ends when one player manages to connect two opposite sides of the board with a continuous path of their pieces. Naively, checking whether the edges of the board are linked requires the expensive step of computing the board’s connected components after each move. However, updating the board’s connected components as a result of placing a single piece can be done very efficiently (a technique used well in the PGX implementation). At compile time, Ludax determines whether a game makes use of a “connection” rule and modifies the step function to iteratively update and track the board’s connected components if so, greatly speeding up later checks. In future extensions, this functionality will be used to accommodate games with atypical or computationally expensive rules without affecting the runtime of existing games.
Figure 2: Average throughput (moves per second) on various exemplar games for Ludax, Ludii, and PGX. The top row of games are implemented in all three frameworks, while the bottom row of games are implemented only in Ludax and Ludii. Speeds for Ludax and PGX are reported for 500 episodes of various batch sizes on a workstation with a single NVIDIA 4090 GPU and 32 CPU cores, while speeds for Ludii are reported for parallel execution on the same workstation across 1, 16, and 32 threads. Error bars are standard deviations calculated over the 500 episodes.
5 Expressive Range
As mentioned above, Ludax currently supports a relatively narrow class of games: two-player, perfect-information board games played by placing, capturing, and flipping pieces. Both Ludii and PGX contain many games that Ludax does not: PGX includes implementations of Backgammon, Chess, Shogi, and Go, along with hidden-information and Atari-style games (though Ludax also supports games that do not appear in PGX), while Ludii’s description language can encode a vast array of board game styles and subgenres. Despite this, Ludax’s description language remains quite expressive. In addition to simple $m-n-k$ line completion games, Ludax supports complex and asymmetric winning conditions (e.g. misère variants, score-based victory), piece capturing and flipping, directional adjacency checks and restrictions, “custodial” mechanics, and games based on connecting arbitrary board regions. Ludax also supports regular rectangular and hexagonal boards of arbitrary sizes, as well as “hexagonal-rectangular” boards (e.g. as used in Hex). These components can then be combined compositionally to form a wide array of unique mechanics and dynamics. In addition, because Ludax is a general description language, implementing a single new game component expands the entire space of games in the framework. While the class of games representable in Ludax may at present be smaller than that of Ludii or other game description languages, it remains expansive.
Figure 3: Performance of reinforcement learning agents trained in the Ludax and PGX implementations of Reversi against the PGX baseline agent. On the left, we plot the average winrate of the learned agents against the baseline over time and across three separate runs. On the right, we plot the average and variance of the winrates. Each run took roughly 3 hours to complete on a workstation with a single A100 GPU.
6 Benchmarking
We benchmark the speed of Ludax on a set of 8 games, 4 of which are also implemented in both Ludii and PGX (allowing for a full comparison) and 4 of which are implemented only in Ludax and Ludii. Again, we emphasize that these 8 games are just exemplars of the class of games which Ludax supports, not an exhaustive list. A full description of each benchmark game is available in the Supplementary Material. We perform each of our benchmarking experiments on a workstation with a single NVIDIA 4090 GPU, 32 CPU cores, and 128GB of memory. In Figure 2 we plot the throughput (in steps per second) under a uniformly random action policy for each game environment against the batch size (log scale on both axes), with the standard deviation of throughputs across episodes as error bars. Ludii supports parallelization via multi-threading: we report throughput on the same workstation when parallelized on 1, 16, and 32 threads. Evaluations for Ludax and PGX were obtained by performing 100 warmup full-game episodes at the specified batch size, followed by measuring the speed over 500 episodes, with each evaluation taking at most a few minutes to complete. Evaluations for Ludii were obtained by running warmup episodes for 10 seconds, followed by measuring the speed over 30 seconds of episodes. [^2] We opted to measure speed for Ludax and PGX using a fixed number of episodes because JAX’s compilation procedure makes it difficult to halt execution after a specific elapsed wall time.
Overall, Ludax achieves speeds that are competitive with state-of-the-art JAX environments. At small batch sizes, its throughput is similar to that of the PGX implementations. At larger batch sizes in more complicated games (i.e. Hex and Reversi), PGX takes a clear edge – though Ludax remains within an order of magnitude of PGX. The comparative “plateauing” of Ludax’s speed at high batch sizes may be due to memory pressure – for instance, Ludax’s implementation of Hex maintains both a board and the connected components for each game state, whereas the PGX implementation cleverly combines both into a single array. This kind of optimization is of course theoretically implementable in Ludax as well, though again we emphasize the desiderata of avoiding game-specific solutions.
Ludax also outspeeds Ludii on 16 and 32 threads across all 8 games, achieving a maximum speedup of between $\sim$3x (Hex) and $\sim$55x (Pente). We note that there are factors that both advantage and disadvantage Ludax in this specific comparison against Ludii. One potential advantage for Ludax is its smaller representation space – implementations of basic mechanics in Ludii support a wider range of optional arguments and board types, with a corresponding increase in computational overhead (though see Section 4 for how this may be avoided). Conversely, Ludii’s ability to use dynamically-sized data structures brings advantages that are particularly beneficial in uniformly random playouts, but would (partially) disappear in playouts using deep reinforcement learning. Specifically, where Ludax samples random actions from categorical probability distribution by masking over the entire board at each step, Ludii selects random actions simply by taking a random index from a list. Ludii also has optimized playout implementations tailored towards the many of the categories of games covered by Ludax Soemers_2022_Optimised , though these optimizations are also more difficult to apply in the context of deep learning.
7 Learned Agents
Finally, we demonstrate the feasibility of training reinforcement learning agents using the Ludax framework. We train our agent on the game Reversi (also known as Othello) using the AlphaZero-style silver2017mastering training script from the PGX library [^3] https://github.com/sotetsuk/pgx/blob/main/examples/alphazero/train.py (used under Apache 2.0 license) (making only slight modifications to accommodate minor differences between the Ludax and PGX APIs). We use the same ResNetV2 kaiming2016resnet network architecture and training hyperparameters as PGX (full details available in the Supplementary Material) and train three separate runs on a single A100 GPU. Each run lasted roughly 57 million frames and took roughly three hours to complete.
We compare the performance of agents trained in the Ludax and PGX environments against the baseline Reversi agent provided by the PGX library in Figure 3. Evaluations were performed by playing two batches of 1024 games (one with the learned agent as the first player and one as the second player), with actions sampled from the normalized output of the policy head at each step. We see that both learned agents achieve remarkably similar performances against the baseline, with little to no differences in learning speed or stability. While a more thorough, tournament-based evaluation would be necessary to properly rank the agents against each other, our objective is to demonstrate the general success of the training procedure and not to definitively defeat the baseline agent. Although the PGX implementation of the Reversi environment is slightly more efficient, this translated into only marginal improvements in overall runtime owing to the shared overhead of network forward passes and weight updates. Like PGX, Ludax offers a familiar API and an efficient set of implementations with which to train learned player agents.
8 Limitations
Generality: As mentioned in Section 5, Ludax currently supports a smaller class of games than other comparable game description languages. While we aim to increase the range of games expressible in Ludax (see below), it will likely never match the full generality of Ludii. As such, other frameworks may be more appropriate for use-cases in which a broad range of games is more important than rapid simulation. Further, Ludax does not support genres other than board games (e.g. video games, card games, …) – we leave the development of hardware accelerated description languages for such domains as an exciting area of future work.
Efficiency: Compared to bespoke JAX implementations of board games (such as in the PGX library), environments in Ludax have worse throughput – especially at larger batch sizes. While we deploy a number of optimizations to help close the efficiency gap when possible (see Section 4), there are ultimately unavoidable trade-offs between speed and generality. For the purpose of training or benchmarking single-task agents on existing games, hard-coded simulators are likely the superior choice. We note, however, that Ludax also provides a way for non-experts to design and benchmark on novel games without having to write any JAX code themselves.
9 Future Work
The most obvious avenue of extension for Ludax is the implementation of additional game mechanics. In particular, we aim to support games based on piece movement (e.g. Amazons), games with multiple piece types (e.g. Checkers), and games with multiple distinct gameplay phases (e.g. Nine-Men’s Morris). Other generalizations, such as support for irregular board shapes, would also help expand the space of games representable in Ludax. In addition, it’s also very likely that the implementation of specific gameplay elements could be further optimized for throughput and / or memory footprint. However, a balance must be struck between efficiency and generality: a less efficient solution which accommodates all valid games under the grammar is ultimately preferable to one which only applies to a subset of games. Lastly, we aim to provide a more robust visual interface for Ludax, both for the purpose of facilitating human-subject research and the potential development of more “human-like” artificial agents which process the game board at the pixel level and select actions spatially.
We are particularly excited about the potential application of Ludax to the study of automated game design (or reward-guided program synthesis more generally Cui_2021_AlphaEvolve surina2025algorithmdiscoveryllmsevolutionary RomeraParedes2024FunSearch ). Systems like GAVEL Todd_2024_GAVEL depend on both a broad representation space and rapid evaluation of novel games. The efficiency of Ludax may make it possible to train a reinforcement learning agent from scratch as part of the inner loop of game evaluation, potentially unlocking a new range of computational features (e.g. learning curves) that correlate with human notions of fun and engagement. Relatedly, Ludax may prove useful to research on human behavior and play. Recent work has explored heuristic-based computational models of human play on simple line completion games zhang2024people , and Ludax offers the possibility to both accelerate computation and broaden the domain to a wider class of games. Finally, Ludax offers an avenue to extend recent research in general game playing (e.g. with large language models schultz2024mastering ) by providing a wide base of efficient game implementations that can in turn be leveraged for tree search algorithms or training world models.
10 Conclusions
We introduce a novel framework for games research that combines the generality of game description languages with the efficiency of modern hardware-accelerated learning environments. Our framework, Ludax, represents a broad class of two-player board games and compiles directly into code in the JAX Python library. Games in Ludax achieve speeds that are competitive with hand-crafted JAX implementations and faster than the widely-used Ludii game description language, and Ludax environments can easily be deployed in existing pipelines for deep reinforcement learning. Our framework helps generalize and accelerate games research, with the potential to unlock entirely new kinds of agents and systems.
Broader Impact
This paper presents a general framework with the goal of advancing reinforcement learning and games research. While there are many potential societal consequences of such work in general, we do not feel that any must be specifically highlighted here.
Appendix A Ludax Grammar
Below we present the complete grammar specification for Ludax, using the syntax of the Lark Python library.
game: "(game" name players equipment rules ")"
// ---Players---
players: "(players" positive_int ")"
// ---Equipment---
equipment: "(equipment" board ")"
board: "(board" (board_square | board_rectangle | board_hexagon | board_hex_rectangle) ")"
board_square: "(square" number ")"
board_rectangle: "(rectangle" number number ")"
board_hexagon: "(hexagon" number ")"
board_hex_rectangle: "(hex_rectangle" number number ")"
// ---Rules---
rules: "(rules" start_rules? play_rules end_rules ")"
// ---Start rules---
start_rules: "(start" start_rule+ ")"
start_rule: start_place
start_place: "(place" player_reference pattern_arg ")"
// ---Play rules---
play_rules: "(play" play_phase+ ")"
play_phase: phase_once_through | phase_repeat
phase_once_through: "(once-through" play_mover_order play_mechanic ")"
phase_repeat: "(repeat" play_mover_order play_mechanic ")"
play_mover_order: "(" player_reference+ ")"
play_mechanic: play_place force_pass?
play_place: "(place" mover_reference? place_destination_constraint place_result_constraint? play_effects?")"
force_pass: "(force_pass" ")"
// ---Constraints---
place_destination_constraint: "(destination" super_mask ")"
place_result_constraint: "(result" super_predicate ")"
// ---Effects---
play_effects: "(effects" play_effect+ ")"
play_effect: effect_capture
| effect_flip
| effect_increment_score
| effect_set_score
effect_capture: "(capture" super_mask mover_reference? increment_score_arg? ")"
effect_flip: "(flip" super_mask mover_reference? ")"
effect_increment_score: "(increment_score" mover_reference function ")"
effect_set_score: "(set_score" mover_reference function ")"
// ---Functions---
function: function_add
| function_connected
| function_constant
| function_count
| function_line
| function_multiply
| function_score
| function_subtract
function_add: "(add" function+ ")"
function_connected: "(connected" multi_mask_arg mover_reference? direction_arg? ")"
function_constant: positive_int
function_count: "(count" super_mask ")"
function_line: "(line" positive_int orientation_arg? exact_arg? ")"
function_multiply: "(multiply" function+ ")"
function_score: "(score" mover_reference ")"
function_subtract: "(subtract" function function ")"
// ---End rules---
end_rules: "(end" end_rule+ ")"
end_rule: "(if" super_predicate end_rule_result ")"
?end_rule_result: result_win | result_lose | result_draw | result_by_score
// -- Result definitions --
result_win: "(" mover_reference "win" ")"
result_lose: "(" mover_reference "lose" ")"
result_draw: "(" "draw" ")"
result_by_score: "(" "by_score" ")"
// -- Mask definitions --
super_mask: mask | super_mask_and | super_mask_or | super_mask_not
super_mask_and: "(and" super_mask+ ")"
super_mask_or: "(or" super_mask+ ")"
super_mask_not: "(not" super_mask ")"
mask: mask_adjacent
| mask_center
| mask_corners
| mask_custodial
| mask_edge
| mask_empty
| mask_occupied
| mask_pattern
| mask_prev_move
mask_adjacent: "(adjacent" super_mask direction_arg? ")"
mask_center: "center"
mask_corners: "corners"
mask_custodial: "(custodial" custodial_length_arg mover_reference? orientation_arg? ")"
mask_edge: "(edge" edge ")"
mask_empty: "empty"
mask_occupied: "occupied" | "(occupied" mover_reference ")"
mask_pattern: "(pattern" dimensions_arg pattern_arg rotate_arg? ")"
mask_prev_move: "(prev_move" mover_reference ")"
// "Multi-masks" are special keywords that are manually split into multiple
// sub-masks at compile time. This is mostly useful for the "connected" function,
// which expects a list of masks to check for connections between
multi_mask: multi_mask_corners
| multi_mask_edges
| multi_mask_edges_no_corners
multi_mask_corners: "corners"
multi_mask_edges: "edges"
multi_mask_edges_no_corners: "edgesNoCorners"
// ---Predicate definitions---
super_predicate: predicate | super_predicate_and | super_predicate_or | super_predicate_not
super_predicate_and: "(and" super_predicate+ ")"
super_predicate_or: "(or" super_predicate+ ")"
super_predicate_not: "(not" super_predicate ")"
predicate: predicate_equals
| predicate_exists
| predicate_full_board
| predicate_function
| predicate_greater_equals
| predicate_less_equals
| predicate_mover_is
| predicate_passed
predicate_equals: "(=" function+ ")"
predicate_exists: "(exists" super_mask ")" // technically equivalent to (>= (count mask) 1)
predicate_full_board: "(" "full_board" ")"
predicate_function: function // special syntax which is equivalent to "(>= function 1)"
predicate_greater_equals: "(>=" function function ")"
predicate_less_equals: "(<=" function function ")"
predicate_mover_is: "(mover_is" player_reference ")"
predicate_passed: "(passed" (mover_reference | BOTH) ")"
// Additional (potentially optional) arguments for predicates
custodial_length_arg: ANY | positive_int
dimensions_arg: "(" positive_int positive_int ")"
direction_arg: "direction:" direction
exact_arg: "exact:" boolean
increment_score_arg: "increment_score:" boolean
multi_mask_arg: multi_mask | "(" super_mask+ ")"
orientation_arg: "orientation:" orientation
pattern_arg: "(" positive_int+ ")"
rotate_arg: "rotate:" boolean
// General-purpose definitions
?number: SIGNED_NUMBER
?positive_int: /[0-9]+/
?boolean: TRUE | FALSE
?edge: TOP | BOTTOM | LEFT | RIGHT | TOP_LEFT | TOP_RIGHT | BOTTOM_LEFT | BOTTOM_RIGHT
?direction: UP | DOWN | LEFT | RIGHT | UP_LEFT | UP_RIGHT | DOWN_LEFT | DOWN_RIGHT | VERTICAL | HORIZONTAL | ORTHOGONAL | DIAGONAL | BACK_DIAGONAL | FORWARD_DIAGONAL | ANY
?orientation: VERTICAL | HORIZONTAL | ORTHOGONAL | DIAGONAL | BACK_DIAGONAL | FORWARD_DIAGONAL | ANY
?player_reference: P1 | P2
?mover_reference: MOVER | OPPONENT
name: STRING
variable_name: /\?[a-z][a-z0-9]*/
id: /[a-zA-Z0-9_]+/
// Constants
TOP: "top"
BOTTOM: "bottom"
UP: "up"
DOWN: "down"
LEFT: "left"
RIGHT: "right"
TOP_LEFT: "top_left"
TOP_RIGHT: "top_right"
BOTTOM_LEFT: "bottom_left"
BOTTOM_RIGHT: "bottom_right"
UP_LEFT: "up_left"
UP_RIGHT: "up_right"
DOWN_LEFT: "down_left"
DOWN_RIGHT: "down_right"
VERTICAL: "vertical"
HORIZONTAL: "horizontal"
ORTHOGONAL: "orthogonal"
DIAGONAL: "diagonal"
BACK_DIAGONAL: "back_diagonal"
FORWARD_DIAGONAL: "forward_diagonal"
ANY: "any"
TRUE: "true"
FALSE: "false"
MOVER: "mover"
OPPONENT: "opponent"
P1: "P1"
P2: "P2"
BOTH: "both"
Appendix B Benchmark Game Descriptions
Below, we present natural language descriptions of the rules for each of the exemplar games analyzed in Section 6.
Tic-Tac-Toe: Players take turns placing a piece into an empty space on a square 3-by-3 board. If a player forms a line of three of their pieces in a row (either vertically, horizontally, or diagonally), they win. If the board is completely full but no lines have been formed, then the game ends in a draw.
Connect Four: Players take turns placing a piece into the top of one of the seven columns on a 6-by-7 board. The piece then “falls” until it rests on either the bottom of the board or another piece. A player can’t place a piece into a column that is already “full.” If a player forms a line of four of their pieces in a row (either vertically, horizontally, or diagonally), they win. If the board is completely full but no lines have been formed, then the game ends in a draw.
Hex: Players take turns placing a piece into an empty space on an 11-by-11 board composed of hexagonal tiles (forming a parallelogram, see visual depiction here). The objective for the first player is to form a continuous path of their pieces that connects the top edge of the board with the bottom edge, while the objective for the second player is to do the same but connect the left and right edges of the board. The first player to achieve their objective wins the game. Because of the geometric properties of the board, it’s not possible for the game to end in a draw.
Reversi: The game takes place on a square 8-by-8 board. To begin, a white piece is placed at positions D4 and E5 and a black piece is placed at positions D5 and E4 (see visual depiction here). Players take turns placing a piece into an empty space such that a line of one or more of the opponent’s pieces are “sandwiched” on either end by the player’s pieces. This configuration is called a “custodial” arrangement of pieces. After placing a piece, any of the opponent’s pieces which are in such a custodial arrangement are flipped and now belong to the player who just moved. It’s possible for a single move to form multiple custodial arrangements in different directions, in which case all of the relevant pieces are flipped. If a player cannot make a legal move, they must pass (and they cannot pass without making a move otherwise). If both players pass, then the game is over. The winner is determined by the player who has the largest number of pieces on the board at the end of the game (in the event of a tie, the game ends in a draw).
Gomoku: Players take turns placing a piece into an empty space on a square 15-by-15 board. If a player forms a line of exactly five of their pieces in a row (either vertically, horizontally, or diagonally), they win. However, forming a line of six or more does not count – the player must have at least one line of exactly five. If the board is completely full but no lines of exactly five have been formed, then the game ends in a draw.
Pente: Players take turns placing a piece into an empty space on a square 19-by-19 board. If a player forms a line of five of their pieces in a row (either vertically, horizontally, or diagonally), they win. In addition, if placing a piece causes a line of exactly two of the opponent’s pieces to be put into a custodial arrangement, the two pieces are captured and removed from a board. Note that placing a piece into a custodial arrangement formed by the opponent does not result in any pieces being captured. A player who captures at least 10 of the opponent’s pieces over the course of the game wins. In the variant of Pente implemented in Ludii and Ludax, the first player must make their first move into the exact center of the board.
Yavalath: Players take turns placing a piece into an empty space on a regular hexagonal board with a diameter of 9 spaces. If a player forms a line of four of their pieces in any direction (either diagonally or horizontally [^4] Ludax assumes a canonical orientation for hexagonal boards in which the diameter stretches from left to right, though it is functionally equivalent to the orientation in which the diameter runs vertically)), they win. However, if a player forms a line of three of their pieces in a row without also forming a line of four, they lose. If the board is completely full but no lines of four or three have been formed, then the game ends in a draw.
Yavalax: To begin, the first player places a piece into an empty space on a square 13-by-13 board. Starting with Player 2, players then take turns placing two pieces into empty spaces on the board. If a player forms at least two distinct lines of four of their pieces in any direction (either vertically, horizontally, or diagonally), they win. However, a player may not place a piece into a space if doing so would form a line of five pieces in any direction or if it would form exactly one line of four pieces in any direction. Note that this restriction applies to a player’s first move of their turn even if they could form a second line of four pieces with their second move of the turn (and thus win). If the board is completely full and neither player has formed at least two distinct lines of four pieces, then the game ends in a draw.
Appendix C Training Hyperparameters
Below we provide the exact training hyperparameters used in the reinforcement learning experiments in Section 7. These are largely copied from the PGX implementation.
  • Model architecture: ResnetV2
  • Number of channels: 128
  • Number of layers: 6
  • Self-play batch size: 1024
  • Self-play simulations: 32
  • Self-play max steps: 256
  • Training batch size: 4096
  • Learning rate: 0.001
  • Evaluation frequency: 5
  • Training iterations: 219
Note that each “iteration” consists of generating play data for 256 steps using the self-play batch size of 1024 (see [22]). We train the model for 219 iterations, which corresponds to $256\times 1024\times 219=57409536$ (or roughly $57$ million) steps in the environment.